# New Multi-messenger Probe of Dark Matter-Nucleon Interactions from Ultra-high Energy Cosmic Ray Acceleration

### v1

@@ III DM Cosmic Ray Interactions
There are two relevant quantities needed for the discussion here. One is the momentum transfer during a CR and DM interaction event. The maximal momentum transfer can be written as [Bringmann:2018cvk, Cappiello:2018hsu, Dent:2019krz, Ciscar-Monsalvatje:2024tvm]
@@
Here we consider the energy transfer in the nucleus’s rest frame, with $Q^{2}=-t$, where $t$ is the Mandelstam variable. Setting $Q^{2}=-(p_{\rm\chi}-p_{\chi}^{\prime})^{2}>0$, we get $Q^{2}=-2m_{\chi}^{2}+2m_{\chi}E_{\chi}^{\prime}=2m_{\chi}T_{\chi}^{\prime}<2m_{\chi}T_{\rm max}$. We now make the conservative requirement that $\sqrt{Q^{2}}\gtrsim 1$ GeV for iron to shatter, far above the binding energy of iron, $\epsilon_{b}\sim 9$ MeV, to ensure a thorough fragmentation. Therefore, requiring $m_{\chi}>1\text{ GeV}^{2}/2T_{\rm max}\simeq 1\text{ GeV}^{2}/2T_{\rm CR}$ implies that for $50$ EeV, we require $m_{\chi}\gtrsim 0.01$ eV. For a more detailed discussion on momentum transfers and cross sections, see [Bardhan:2022bdg, Zhang:2025rqh].
@@
Note that at the energies discussed here, we are beyond the regime of elastic scattering. For instance, for the scattering of a $T_{\mathrm{iron}}=100$ TeV iron CR on a $m_{\chi}=0.1$ MeV DM target, the average momentum transfer is $\langle{\Delta T}\rangle\sim 4.5$ GeV, demonstrating that we are safely within the regime of deep inelastic scattering.
@@ IV Iron Nuclei Fragmentation
Here we focus on the possibility of accelerating iron to ultra-high energies near the SMBH and/or the inner jet, assuming a background dark-matter distribution with couplings to protons. See [Rieger:2019wgt, Rieger:2022qhs] for reviews of possible models. Specifically, we ask the question: Would a coupling to DM shatter the iron nuclei before it can accelerate to observed ultra-high energies?
@@ IV.1 Accelerating within an AGN
Assuming UHECR near the GZK cut-off ($E_{\mathrm{CR}}\sim 50\;\mathrm{EeV}$) are all accelerated within AGNs, we can set a basic requirement on the interaction length for CR-DM interactions, $\lambda_{\mathrm{CR},\chi}$. The interaction length $\lambda_{\mathrm{CR},\chi}$ needs to be larger than the acceleration distance $l_{\mathrm{acc}}$ of the CR. Otherwise, the CR (at these extreme energies assumed to be iron) would shatter and could not be accelerated to the necessary energies. Conservatively, we can set $l_{\mathrm{acc}}$ to the size of the AGN. Current observations set the size of AGNs to be $\lesssim 10^{-3}\;\mathrm{pc}$ [2022MNRAS.511.3005J, 2022ApJ...929...19G]. Note here we define the size of the AGN to be the size of the accretion disk of the central SMBH.
Thus, assuming CR are accelerated within the accretion disk of the SMBH, we can set a conservative bound by setting $\lambda_{\chi,\mathrm{CR}}\geq 10^{-3}\;\mathrm{pc}$. Using the previously defined DM-density profiles (at $r=10^{-3}\;\mathrm{pc}$), we can then set a generic bound on $\sigma_{\chi,\mathrm{CR}}$
@@
While these bounds are negligible for DM masses above 1 keV compared to current constraints [LZ:2025iaw, Nadler:2019zrb], for lower masses they begin to dominate over current direct detection constraints.
@@ IV.2 Accelerating within its trapped region
We can refine the estimate in Equation 6 by observing that if a particle escapes its acceleration region, it no longer accumulates energy. i.e., the Larmor radius $r_{L}=E/(ZeB)$ needs to be smaller than the characteristic size of the source, $L$. Here, $B$ is the magnetic field strength within the source, $e$ the unit charge, and $Z$ the charge number of the CR nucleus. This is called the Hillas Criterion [Hillas:1984ijl]. As discussed in Refs. [Hillas:1984ijl, Ptitsyna:2008zs, Kotera:2011cp, Rieger:2022qhs] this is a necessary but not sufficient constraint on CR acceleration. Thus, we can treat it as a conservative scale. The criterion can be written as
$E_{\mathrm{CR}}<10^{20}Z\left(\frac{B}{10\,\mu\mathrm{G}}\right)\left(\frac{L}{10\;\mathrm{kpc}}\right)\;\mathrm{eV}.$ (7)
Note that this criterion can be improved by introducing flow velocities within the object [Ptitsyna:2008zs, Kotera:2011cp, Bell:2013vxa, Rieger:2022qhs]. Here, we change the criterion to the following: The particle’s interaction length, $\lambda_{\chi,\mathrm{CR}}$, needs to be larger than the Larmor-radius, $r_{L}$. This is equivalent to requiring that the CR does not interact before escaping the acceleration site. Following [Rieger:2022qhs] we define $r_{L}$ as
$r_{L}=1.08\frac{\frac{E}{\mathrm{PeV}}}{ZB_{\mu\mathrm{G}}}\mathrm{pc},$ (8)
where $B_{\mu\mathrm{G}}$ is the magnetic field in micro-Gauss. To be conservative, we impose our criterion using the smallest Larmor radius imaginable at the source, namely that obtained with the magnetic field strength at the horizon of the SMBH. From observations, the magnetic field strength at the SMBH’s horizon was estimated to be $B\sim 10^{4}\;\mathrm{G}$ [2009A&A...507..171S, Baczko:2016opl, Piotrovich:2020ooz]. Thus, the Larmor radius for iron nuclei at the GZK cut-off in an AGN is
@@
Another reason why this should be seen as a conservative bound is that it assumes the acceleration efficiency to be maximal, $\eta=1$. As discussed in [Aharonian:2002we], $\eta$ within these objects can be as low as $\eta\geq 0.03$. Note, other authors (e.g. [2009JCAP...11..009L]) describe this efficiency in terms of the acceleration time, $t_{\mathrm{acc}}$. In general, it can be written as
@@
with $\mathcal{A}$ being a multiplier of the Larmor time $t_{\mathrm{L}}$. In general, $\mathcal{A}\gg 1$ and in the specific case $\mathcal{A}\approx 1$, this corresponds to a maximally efficient acceleration process in a semi-relativistic scenario [2009JCAP...11..009L] (the assumption we are using here). In this scenario $t_{\mathrm{acc}}\sim t_{\mathrm{scattering}}$, where $t_{\mathrm{scattering}}$ is the scattering time-scale. We now apply these criteria and the resulting bounds are
@@
improving the previous estimate. Note that the improvements are far smaller for an assumed NFW profile.
These bounds can now be further improved by assuming all CR are accelerated within an NGC1068-like galaxy, instead of a MW-like one. Plugging in the numbers from [Herrera:2023nww] (and $R_{\mathrm{emission}}\leq 10^{-4}\;\mathrm{pc}$) we get a significant improvement of the bounds
@@ IV.3 Can the Cosmic Ray escape the galaxy?
$\sigma_{\chi,\mathrm{CR}}\leq\left\{\begin{array}[]{ll}3\times 10^{-24}\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)\;\mathrm{cm^{2}}&\mathrm{Burkert,}\\[6.0pt] 3\times 10^{-26}\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)\;\mathrm{cm^{2}}&\mathrm{NFW,}\\[6.0pt] 9\times 10^{-31}\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)\;\mathrm{cm^{2}}&\mathrm{MW-Spike,}\\[6.0pt] 1\times 10^{-33}\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)\;\mathrm{cm^{2}}&\mathrm{NGC-Spike}.\end{array}\right$. (13)
This can be understood by comparing these numbers to the expectations from the GZK cut-off. For the GZK limit the photon-CR cross-section $\sigma_{\gamma p}\sim 3\times 10^{-28}\;\mathrm{cm^{2}}$ is the relevant quantity [Arkhipov:2006vy]. The CMB denstiy per $\mathrm{cm}^{3}$ is approximately 411 photons. This implies an interaction length of $\sim 10$ Mpc. Here the high DM densities scale the resulting cross-section for DM to lower values. For instance, the average density in a DM spike for a Milky-Way-like galaxy between $2\times 10^{-5}\;\mathrm{pc}$ and $1$ pc is $\sim 7\times 10^{9}\;\mathrm{GeV/cm^{3}}$.
Note, that the bound set by escape is stronger than the bound due to acceleration, mainly due to our conservative assumptions on the Larmor radius and acceleration time. Had we used smaller magnetic field strengths, the constraint from the trapped region would be stronger. For example, setting $B=100\;\mathrm{G}$, results in $r_{L}\sim 4\times 10^{-5}\;\mathrm{pc}$, similar in size to the assumed location of acceleration. This alone would yield a bound due to the acceleration criteria similar to the escape shown in Equation 13. If the acceleration were not 100% efficient, this would only further improve.
@@ IV.4 A non-conservative estimate
While in the previous sections, we used conservative estimates to derive bounds on DM-CR interactions, here we relax them: In NGC 1068 CRs can be accelerated as close as $\sim 10^{-5}$ pc from the central SMBH [Herrera:2023nww]. Additionally, as discussed previously, the efficiency of acceleration could be as low as $\eta=0.03$ [Aharonian:2002we] (see [Yang:2024bsf] for even lower efficiencies). With these numbers $r_{L}\approx 1.3\times 10^{-5}$ pc and $\rho_{\chi}^{\rm NGC}(r=10^{-5}\;\mathrm{pc})\approx 5\times 10^{18}\;\mathrm{GeV/cm^{3}}$. In this scenario, the acceleration bound is the dominant contribution and the resulting constraint on the cross-section is $\sigma_{\chi,{\rm CR}}\leq 8\times 10^{-35}\;\frac{m_{\chi}}{\mathrm{GeV}}\;\mathrm{cm^{2}}$.
@@ IV.5 Baryonic Feedback
Here we provide a rough estimate of baryonic feedback caused by the $\chi$-$p$ coupling on the formation of the spike. Baryonic feedback is relevant when DM has scattered at least once during the lifetime of the galactic DM halo, $t_{\mathrm{age}}$. In that case, in order to avoid baryonic feedback, we can require that
@@
For illustration, we can use the age of the MW halo, $t_{\mathrm{age}}\sim 10\times 10^{9}$ years, the stellar mass density near Sagittarius A, $\rho_{\odot}\sim 3\times 10^{7}\;M_{\odot}\;\mathrm{pc^{-3}}$, and converting this criterion to a bound on the cross section, we obtain the approximate requirement that
$\sigma_{\chi p}\leq 10^{-34}\;\mathrm{cm^{2}}$ (15)
to avoid feedback. The reasoning above assumes an isothermal core formation [BetancourtKamenetskaia:2025ivl] (see also [Spergel:1999mh, Kaplinghat:2015aga]).
For larger cross sections (e.g., for most of the parameter space we target in the Milky-Way scenario), we assume the DM density profile follows the strong stellar heating scenario shown in Figure 2 and calculated in [Balaji:2023hmy]. Figure 4 shows the resulting bounds including baryonic feedback. For the Milky Way scenario, this applies for all DM masses, while for the NGC 1068 case, masses $\geq 100\;\mathrm{MeV}$ are affected, which appears in our curves as a kink at about $m_{\chi}\sim 100$ MeV. We note that NGC1068 is significantly younger source, so the bound in Equation 15 will weaken accordingly.
Figure 3: The resulting bounds using the NFW and Spike profiles for the Milky-Way (marked MW) and assuming NGC1068-like population. These bounds are constructed under the assumption of a GZK-like process. The experimental bounds are from LZ [LZ:2025iaw]. Direct detection bounds adopted from [Billard:2021uyg] with bounds from CRESST [CRESST:2017ues, CRESST:2022dtl], DAMIC [DAMIC:2020cut], DarkSide [DarkSide:2018bpj, DarkSide:2018kuk], and XENON1T[XENON:2018voc, XENON:2019gfn]. BBN [Sabti:2019mhn, Krnjaic:2019dzc, Sabti:2021reh].
Figure 3 shows the resulting bounds for spiked profiles, compared to current constraints.
@@ IV.6 A comment on TXS 0506+056
TXS 0506+056 was observed to be a source of neutrinos by Icecube [IceCube:2018cha]. Based on the neutrino and Gamma-Ray observations [2017ATel10791....1T, MAGIC:2018sak], a few models predict that the main acceleration region of CR is $\mathcal{O}$(pc) away from the central SMBH [deGouveiaDalPino:2024txs, Kun:2018txs, Caproni:2022txs], well outside of the possible DM spike. On the other hand, models exist where the production of neutrinos (thus the acceleration of CRs) happens between $4\times 10^{-4}-0.3$ pc [Xue:2019txw, Xue:2020kuw, Yang:2024bsf]. Based on this, the spike profile has been used to constrain DM-neutrino interactions in those scenarios [Zapata:2025huq].
Setting the mass of the central SMBH to be $3\times 10^{8}\;M_{\odot}$ [Padovani:2019xcv], the dark matter density will be between $10^{10}-10^{15}$ GeV/cm3 for the distances of interest. This is far. Assuming all high-energy (iron) CRs are accelerated within TXS-like objects, this would lead to a bound of $\sigma_{\chi,CR}\lesssim 10^{-33}\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)$ cm2, when considering the CR should be able to escape the spike without interacting. This is similar to the NGC bound and is mainly due to the similar column densities between the two when considering escaping CRs.
Figure 4: The bounds for strongly interacting DM compared to current constraints. Current constraints are from XQC [Erickcek:2007jv], CMB [Gluscevic:2017ywp], Upscattered DM (LZ-Boosted [LZ:2025iaw] and PROSPECT [PROSPECT:2021awi]), EDELWEISS [EDELWEISS:2019vjv], CRESST 2017 (surface) [CRESST:2017ues, Emken:2018run], XENON1T [XENON:2017vdw, XENON:2018voc, XENON:2019gfn], CRESST-III [CRESST:2019jnq, CRESST:2019axx], DAMIC [DAMIC:2020cut], DarkSide [DarkSide:2018bpj, DarkSide:2018kuk], and XENON1T[XENON:2018voc, XENON:2019gfn]. BBN [Sabti:2019mhn, Krnjaic:2019dzc, Sabti:2021reh]. Milky-Way Sattelites [Nadler:2019zrb], Lyman-$\alpha$ [Rogers:2021byl], and re-interpretations of existing data [Emken:2019tni].

### v2

@@ III DM Cosmic Ray Interactions
There are two relevant quantities needed for the discussion here. One is the momentum transfer during a CR and DM interaction event. The maximal momentum transfer can be written as [77, 56, 58, 71, 61]
@@
Here we consider the energy transfer in the nucleus’s rest frame, with $Q^{2}=-t$, where $t$ is the Mandelstam variable. Setting $Q^{2}=-(p_{\rm\chi}-p_{\chi}^{\prime})^{2}>0$, we get $Q^{2}=-2m_{\chi}^{2}+2m_{\chi}E_{\chi}^{\prime}=2m_{\chi}T_{\chi}^{\prime}<2m_{\chi}T_{\rm max}$. We now make the conservative requirement that $\sqrt{Q^{2}}\gtrsim 1$ GeV for iron to shatter, far above the binding energy of iron, $\epsilon_{b}\sim 9$ MeV, to ensure a thorough fragmentation. Therefore, requiring $m_{\chi}>1\text{ GeV}^{2}/2T_{\rm max}\simeq 1\text{ GeV}^{2}/2T_{\rm CR}$ implies that for $50$ EeV, we require $m_{\chi}\gtrsim 0.01$ eV. For a more detailed discussion on momentum transfers and cross sections, see Refs. [41, 175].
@@
Note that at the energies discussed here, we are beyond the regime of elastic scattering. For instance, for the scattering of a $T_{\mathrm{iron}}=100$ TeV iron nucleus on a $m_{\chi}=0.1$ MeV DM target, the average momentum transfer is $\langle{\Delta T}\rangle\sim 4.5$ GeV, demonstrating that we are safely within the regime of deep inelastic scattering.
@@ IV Iron Nuclei Fragmentation
Here we focus on the possibility of accelerating iron to ultra-high energies near the SMBH and/or the inner jet, assuming a background DM distribution with couplings to protons. See Refs. [146, 147] for reviews of possible models. Specifically, we ask the question: Would a coupling to DM shatter the iron nucleus before it can accelerate to observed ultra-high energies?
@@ IV.1 Accelerating within an AGN
Assuming that UHECRs near the GZK cut-off ($E_{\mathrm{CR}}\sim 50\;\mathrm{EeV}$) are all accelerated within AGNs, we can set a basic requirement on the interaction length for CR-DM interactions, $\lambda_{\mathrm{CR},\chi}$. The interaction length $\lambda_{\mathrm{CR},\chi}$ needs to be larger than the acceleration distance $l_{\mathrm{acc}}$ of the CR. Otherwise, the CR (at these extreme energies assumed to be iron) would shatter and could not be accelerated to the necessary energies. Conservatively, we can set $l_{\mathrm{acc}}$ to the size of the AGN. Current observations set the size of AGNs to be $\lesssim 10^{-3}\;\mathrm{pc}$ [111, 100]. Note here we define the size of the AGN to be the size of the accretion disk of the central SMBH.
Thus, assuming CRs are accelerated within the accretion disk of the SMBH, we can set a conservative bound by setting $\lambda_{\chi,\mathrm{CR}}\geq 10^{-3}\;\mathrm{pc}$. Using the previously defined DM-density profiles (at $r=10^{-3}\;\mathrm{pc}$), we can then set a generic bound on $\sigma_{\chi,\mathrm{CR}}$
@@
While these bounds are negligible for DM masses above 1 keV compared to current constraints [2, 133], for lower masses they begin to dominate over current direct detection constraints.
@@ IV.2 Accelerating within its trapped region
We can refine the estimate in Equation 6 by observing that if a particle escapes its acceleration region, it no longer accumulates energy. In other words, the Larmor radius $r_{L}=E/(ZeB)$ needs to be smaller than the characteristic size of the source, $L$. Here, $B$ is the magnetic field strength within the source, $e$ the unit electric charge, and $Z$ the atomic number of the CR nucleus. This is called the Hillas Criterion [104]. As discussed in Refs. [104, 143, 117, 147] this is a necessary but not sufficient constraint on CR acceleration. Thus, we can treat it as a conservative scale. The criterion can be written as
$E_{\mathrm{CR}}<(10^{20}\;\mathrm{eV})Z\left(\frac{B}{10\,\mu\mathrm{G}}\right)\left(\frac{L}{10\;\mathrm{kpc}}\right).$ (7)
Note that this criterion can be improved by introducing flow velocities within the object [143, 117, 43, 147]. Here, we change the criterion to the following: The particle’s interaction length, $\lambda_{\chi,\mathrm{CR}}$, needs to be larger than the Larmor radius, $r_{L}$. This is equivalent to requiring that the CR does not interact before escaping the acceleration site. Following Ref. [147], we define $r_{L}$ as
$r_{L}=\left(\frac{1.08\;\mathrm{pc}}{Z}\right)\left(\frac{E}{1\;\mathrm{PeV}}\right)\left(\frac{1\;\mu G}{B}\right)\,.$ (8)
To be conservative, we impose our criterion using the smallest Larmor radius imaginable at the source, namely that obtained with the magnetic field strength at the horizon of the SMBH. Additionally, we set the acceleration to be maximally efficient, such as in magnetic reconnection [85, 114], meaning the nucleon can accelerate fully within one $r_{L}$.
From observations, the magnetic field strength at the SMBH’s horizon was estimated to be $B\sim 10^{4}\;\mathrm{G}$ [157, 38, 140]. Thus, the Larmor radius for iron nuclei at the GZK cut-off in an AGN is
@@
Another reason why this should be seen as a conservative bound is that it assumes the acceleration efficiency to be maximal, $\eta=1$. As discussed in Ref. [22], $\eta$ within these objects can be as low as $\eta\geq 0.03$. Note, other authors (e.g., Ref. [121]) describe this efficiency in terms of the acceleration time, $t_{\mathrm{acc}}$. In general, it can be written as
@@
with $\mathcal{A}$ being a multiplier of the Larmor time $t_{\mathrm{L}}$. In general, $\mathcal{A}\gg 1$ and in the specific case $\mathcal{A}\approx 1$, this corresponds to a maximally efficient acceleration process as seen in magnetic reconnection [121, 85, 114] (the assumption we are using here). In this scenario $t_{\mathrm{acc}}\sim t_{\mathrm{scattering}}$, where $t_{\mathrm{scattering}}$ is the scattering time-scale. We now apply these criteria and the resulting bounds are
@@
improving the previous estimate for the spike, but reducing it for an NFW profile.
These bounds can now be further improved by assuming all CR are accelerated within an NGC 1068-like galaxy, instead of a MW-like one. Plugging in the numbers from Ref. [102] (and using the assumed emission region $\leq 10^{-4}\;\mathrm{pc}$) we get a significant improvement of the bounds to
@@ IV.3 Can the Cosmic Ray escape the galaxy?
$\sigma_{\chi,\mathrm{CR}}\leq\left\{\begin{array}[]{ll}3\times 10^{-24}\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)\;\mathrm{cm^{2}}&\mathrm{Burkert,}\\[6.0pt] 3\times 10^{-26}\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)\;\mathrm{cm^{2}}&\mathrm{NFW,}\\[6.0pt] 9\times 10^{-31}\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)\;\mathrm{cm^{2}}&\mathrm{MW~Spike,}\\[6.0pt] 1\times 10^{-33}\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)\;\mathrm{cm^{2}}&\mathrm{NGC~Spike}.\end{array}\right$. (13)
[…]
Note that the bound set by escape is stronger than the bound due to acceleration, mainly due to our conservative assumptions on the Larmor radius and acceleration time. Had we used smaller magnetic field strengths, the constraint from the trapped region would be stronger. For example, setting $B=100\;\mathrm{G}$, results in $r_{L}\sim 4\times 10^{-5}\;\mathrm{pc}$, similar in size to the assumed location of acceleration. This alone would yield a bound due to the acceleration criteria similar to the escape shown in Equation 13. If the acceleration were not 100% efficient, this would only further improve.
@@ IV.4 A non-conservative estimate
While in the previous sections, we used conservative estimates to derive bounds on DM-CR interactions, here we relax them: In NGC 1068-like galaxies, CRs can be accelerated as close as $\sim 10^{-5}$ pc from the central SMBH [102]. Additionally, as discussed previously, the efficiency of acceleration could be as low as $\eta=0.03$ [22] (see Ref. [171] for even lower efficiencies). With these numbers $r_{L}\approx 1.3\times 10^{-5}$ pc and $\rho_{\chi}^{\rm NGC}(r=10^{-5}\;\mathrm{pc})\approx 5\times 10^{18}\;\mathrm{GeV/cm^{3}}$. In this scenario, the acceleration bound is the dominant contribution and the resulting constraint on the cross-section is $\sigma_{\chi,{\rm CR}}\leq 8\times 10^{-35}\;\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)\;\mathrm{cm^{2}}$.
@@ IV.5 Baryonic Feedback
Here we provide a rough estimate of baryonic feedback caused by the $\chi$-$p$ coupling on the formation of the spike itself. Baryonic feedback is relevant when DM has scattered at least once during the lifetime of the galactic DM halo, $t_{\mathrm{age}}$. In that case, in order to avoid baryonic feedback, we can require that
@@
For illustration, we can use the age of the MW halo, $t_{\mathrm{age}}\sim 10\;\mathrm{Gyr}\sim 3\times 10^{17}$ seconds. For the baryon (proton) number density, $n_{p}$, we consider the stellar mass density, within the central parsec $\rho_{\star}\sim 10^{6}$$10^{7}\,M_{\odot}/\mathrm{pc}^{3}$ ($\sim 4\times 10^{7}$$4\times 10^{8}\,\mathrm{GeV/cm^{3}}$) [90, 153, 154], the gas density in the inner parsec $n\sim 10$$100/\mathrm{cm}^{3}\sim 10$$100\,\mathrm{GeV/cm^{3}}$ [39], and the central molecular zone $n\sim 10^{7}$$10^{8}/\mathrm{cm}^{3}\sim 10^{7}$$10^{8}\,\mathrm{GeV/cm^{3}}$ [130, 90]. For the velocity, we use $v_{p}\leq 0.1c\simeq 3\times 10^{9}\;\mathrm{cm/s}$, which is above the observed velocity of stars close to Sagittarius A [139]. From this we obtain the approximate requirement that
$\sigma_{\chi p}\leq 10^{-35}\;\mathrm{cm^{2}}$ (15)
to avoid feedback. The reasoning above assumes an isothermal core formation [50] (see also Refs. [159, 113]).
For larger cross sections (e.g., for most of the parameter space we target in the MW scenario), we assume the DM density profile follows the strong stellar heating scenario shown in Figure 2 and calculated in Ref. [40]. Figure 3 shows the resulting bounds including baryonic feedback. For the MW scenario, this applies for all DM masses, while for the NGC 1068 case, masses $\geq 10\;\mathrm{MeV}$ are affected, which appears in our curves as a kink at about $m_{\chi}\sim 10$ MeV. We note that NGC 1068 is a significantly younger source, so the corresponding bound in Equation 15 will weaken accordingly.
Figure 3: The resulting bounds using the NFW and spiked profiles for the MW and assuming NGC 1068-like population. These bounds are constructed under the assumption of a GZK-like process. The experimental bounds are from LZ [2] CR-boosted DM search and the direct detection bounds adopted from Ref. [53] with data from CRESST [30, 31], DAMIC [21], DarkSide [19, 18], and XENON1T[34, 35]. The vertical shaded region is the BBN bound [149, 118, 150].
Figure 3 shows the resulting bounds for spiked profiles, compared to current constraints from CR-boosted DM search [2], direct detection searches [53] and BBN [149, 118, 150].
Figure 4 shows a summary of all relevant experimental constraints, compared to our bounds which rule out new parameter space between in the DM mass region $m_{\chi}\in[3,30]$ MeV for the NGC 1068 case with a spiked profile.
@@ IV.6 A comment on TXS 0506+056
TXS 0506+056 was observed to be a source of neutrinos by Icecube [3]. Based on the neutrino and Gamma-Ray observations [160, 32], a few models predict that the main acceleration region of CR is $\mathcal{O}$(pc) away from the central SMBH [66, 119, 59], well outside of the possible DM spike. Specifically, in Ref. [86] the authors show that a coronal origin (close to the SMBH) is disfavored by X-ray estimations.
On the other hand, models exist where the production of neutrinos (thus the acceleration of CRs) happens between $4\times 10^{-4}-0.3$ pc [168, 169, 171]. Based on this, the spiked profile has been used to constrain DM-neutrino interactions in those scenarios [62, 83, 172].
Setting the mass of the central SMBH to be $3\times 10^{8}\;M_{\odot}$ [137], the DM density will be between $10^{10}-10^{15}$ GeV/cm3 for the distances of interest. Assuming all high-energy (iron) CRs are accelerated within TXS-like objects, this would lead to a bound of $\sigma_{\chi,{\rm CR}}\lesssim 10^{-33}\left(\frac{m_{\chi}}{\mathrm{GeV}}\right)$ cm2, when considering that CR should be able to escape the spike without interacting. This is similar to the NGC bound and is mainly due to the similar column densities between the two when considering escaping CRs.
Figure 4: Summary of our bounds on DM-proton cross section (lines with upward arrows) compared to current constraints from XQC [80], CMB [94], upscattered DM (LZ-boosted [2] and PROSPECT [29]), EDELWEISS [37], CRESST 2017 (surface) [30, 79], XENON1T [33, 34, 35], CRESST-III [11, 10], DAMIC [21], and DarkSide [19, 18], BBN [149, 118, 150], MW Sattelites [133], Lyman-$\alpha$ [148], and re-interpretations of existing data [78].